OpenStudy (anonymous):

How do I use this equation to find the radius of the circle it represents? x^2 + y^2 + 4x + 8y - 10 = 0 I know it's in general circle form. I'm not good with the general form at all, I'm better with the standard form.

OpenStudy (anonymous):

I graphed the equation on a graphing calculator and found that the radius is about 5.5, or so it looks that way... but I figure I have to learn how to actually convert general form to standard form anyway.

OpenStudy (shamil98):

Complete the squares .

OpenStudy (anonymous):

How do I do that.

OpenStudy (shamil98):

x^2 + y^2 + 4x + 8y - 10 = 0 (x^2 + 4x + 4) + (y^2 + 8y + 16) - 10 - 16 - 4 = 0

OpenStudy (anonymous):

complete squares and on right hand side of = you have to complete squares, u will get answer

OpenStudy (shamil98):

(x+2)^2 + (y+4)^2 - 30 = 0 (x+2)^2 + (y+4)^2 = 30

OpenStudy (anonymous):

complete the square

OpenStudy (anonymous):

Oh god this is really confusing lol but I think I'm getting it

OpenStudy (shamil98):

(x-h)^2 + (y-k)^2 = r^2 (h,k) --> (-2,-4) r = √30

OpenStudy (anonymous):

ohhh okay i see what you did there, I get it now!

OpenStudy (anonymous):

I'm gonna have to try it on some other equations, and if I have any problems, well, I'll post another question, lol. Thank you so much!